# Disentangling Vortex-driven Dynamics in the Solar Atmosphere Using Information Theory

### v1

@@ 2.3.1 Coupling analysis using MI
Because MHD systems are inherently nonlinear, traditional correlation-based measures may miss subtle, but important dependencies. MI, however, is sensitive to the full spectrum of possible interactions. To compare how strongly different pairs of variables are connected, or how this connection changes across heights or times, MI is often normalized. This places all variable comparisons on a common scale, making it possible to assess the relative strength of interactions in a consistent manner. Normalization can be done by dividing by the joint entropy, or by the geometric mean of the individual entropies. The result is a number between 0 and 1, where 0 means no shared information, and 1 means complete dependence. For example, a normalized MI value of 0.1 means that the two MHD variables share 10% of their total uncertainty. Put differently, knowing the state of one variable reduces uncertainty about the other by that fraction. In practical terms, a normalized MI above 0.1 indicates a nontrivial statistical dependency. For MHD, this can reveal meaningful coupling between physical processes, such as energy transfer across fields, cross-scale interactions in turbulence, or correlations between magnetic and velocity structures that would not necessarily be visible in linear measures. Thus, even a seemingly “small” MI value may signify deep physical connections in the dynamics of the plasma.
For two random variables and their MI, , is defined as
@@
Before computing the NMI, each variable was rescaled to a fixed range, either or , depending on its physical bounds. Joint histograms were then constructed using 64 uniformly spaced bins across the normalized range. This fixed binning approach provides enough resolution to capture nonlinear dependencies while maintaining numerical stability in the mutual information estimates. Normalizing and binning in this way ensures consistency across variables and time steps, allowing for reliable comparisons and minimizing sensitivity to outliers or local fluctuations in the data.
To assess the statistical reliability of the observed information-theoretic coupling, we performed pixel-shuffle surrogate tests on each time–height window. For each driver–target pair, we generated spatial surrogates by randomly permuting pixel positions within the vortex mask, which preserves marginal distributions while destroying spatial structure. For MI, we computed a null distribution and defined the -value as the fraction of surrogates with values greater than or equal to the observed coupling. Across the entire analysed region and time interval, observed values exceeded all surrogate samples, yielding -values in nearly all pixels. To further quantify the strength of the signal relative to surrogate variability, we computed the effect size as the absolute difference between the observed and mean surrogate values, and defined the signal-to-noise ratio (SNR) as . The resulting SNR maps show extended regions with values exceeding 5, and effect sizes well above the surrogate standard deviation. These metrics demonstrate that the observed couplings are not only statistically significant but also meaningfully stronger than surrogate-induced fluctuations, providing robust evidence of structured dependencies in the analysed regions.
@@ Results
To characterize the temporal and vertical evolution of physical conditions within the vortex, we computed time–height maps of spatially averaged quantities inside the tracked the magnetic flux tube core. By analyzing these averaged fields across the different vortex phases, before onset, development, and decay, we identify the signatures of dynamic coupling and energy concentration associated with vortex-driven activity. Figure 3 shows the temporal and variation with height of the vorticity of horizontal velocity (),vertical velocity () and the base-10 logarithm of the spatially averaged values for each variable within the vortex core. The coloured stars indicate key stages in the vortex evolution as determined by the -method (Graftieaux et al., 2001; Giagkiozis et al., 2018; Yuan et al., 2023), using the condition in the magnetic flux tube mask to define coherent rotation. We choose a lower value as initially the velocity field is not as twisted, leading to a lower . Blue stars mark the onset of the main vortex at each height, while orange stars indicate the end of that initial coherent structure. As the main vortex decays, it fragments into smaller secondary vortices that persist for a time before fully dissipating; their presence is marked by green stars. The vortex appears to form between the high photosphere (approximately 0.45 Mm) and the low chromosphere (around 1.0 Mm), after which it expands both upward into the corona and downward toward the solar surface. In the upper atmosphere, the vortex decays into multiple smaller structures, whereas in the lower layers it disappears more abruptly. In particular, smaller vortices are more persistent around 2 Mm, indicating that vortex evolution may be stratified with height, with different lifetimes and dynamics depending on atmospheric depth.
@@
The onset of the main vortex leaves identifiable signatures across all plasma variables, particularly in the chromosphere and transition region (heights above 1.0 Mm). In the upper atmosphere, a reversal in the sign of the horizontal vorticity is observed around the time of vortex formation. This indicates a transition from small-scale, network-like flows exhibiting clockwise rotation into a dominant large-scale vortex rotating in the counterclockwise direction. The positive vorticity associated with the large-scale structure persists through most of the vortex’s lifetime but gradually weakens as the coherent rotation breaks down. During the decay phase, the vortex transitions into smaller, fragmented swirling motions. This evolution reflects a redistribution of angular momentum and suggests a transfer of energy from organized, large-scale flows into localized, potentially dissipative motions.
@@
One of the most prominent features is a sharp and sustained drop in plasma-, which falls below unity shortly after vortex formation and remains low (log) as long as coherent vortical structures persist. This behaviour marks a transition from a gas-pressure-dominated to a magnetically dominated regime. Simultaneously, both the kinetic and magnetic energy increase by over an order of magnitude, peaking near vortex’s half-life. A local minimum in the vertical magnetic field coincides with the onset of the main vortex (blue stars). This may reflect either a redistribution of magnetic flux by the swirling plasma motions, or the conversion of vertical field into horizontal components as the vortex twists and shears the local field. The subsequent increase in indicates magnetic flux accumulation within the vortex core as it matures.
@@
The formation of the vortex is accompanied by a coherent upflow of material, which transports denser plasma from lower atmospheric layers into the vortex core. This upflow results in a localized enhancement of mass density and a simultaneous drop in temperature, consistent with adiabatic expansion and the lifting of a cooler, denser plasma. Throughout the lifetime of the vortex, upward velocities dominate within the core region, continuously transporting plasma from lower layers and sustaining the elevated density. This persistent upflow plays a crucial role in maintaining the mass loading of the vortex structure. Only toward the end of the vortex’s life does this trend reverse, with downflows gradually becoming more prominent. This culminates in a strong downward jet along the vortex axis during its final decay phase, particularly in the upper atmospheric layers. The gas pressure initially increases at heights above 1.5 Mm as the upflow develops, but subsequently drops as the vortex matures. Following the initial cooling phase, the plasma within the vortex core is gradually heated, reaching temperatures up to an order of magnitude higher than the onset vortex state.
@@ 2.4 Tracking Plasma Coupling and Coherence via Information Theory
The profiles presented in Figure 3 characterize the background evolution of the vortex core and provide a reference context to interpret subsequent information-theoretic and structural analyzes. We computed SE to quantify the local spatial variability of key physical quantities within the tracked magnetic flux tube where the solar tornado is developed. Figure 4 shows the temporal and vertical evolution of the SE for the same variables as Figure 3, except for plasma-. The entropy was computed within the dynamically tracked vortex core, and each panel presents the entropy values, revealing how the spatial organization and complexity of the plasma change over time and height. In practice, entropy values near – correspond to relatively ordered fields with low spatial variability, while intermediate values (–) indicate partially structured states where coherent features coexist with moderate fluctuations. Higher values approaching reflect strongly intermittent fields, likely associated with turbulence, sharp spatial gradients, or enhanced small-scale variability. The entropy was estimated from the empirical probability distributions at each height and time step using a fixed binning scheme of 64 bins.
@@
The onset and development of the main vortex are marked by significant changes in the spatial organization of magnetic fields, as captured by entropy diagnostics. During the formation phase, the entropy of magnetic energy () rises in the lower atmosphere, indicating increasing spatial complexity and localized structuring of the field. This is accompanied by a divergence between the entropy of and that of magnetic shear (), consistent with the formation of current sheets rather than uniform shear enhancement. While becomes more intermittent due to twisting and compression by the vortex, magnetic shear entropy remains vertically patchy and fragmented, suggesting that complexity arises from thin, spatially confined features rather than widespread braiding. Similarly, the entropy of the vertical field component () increases during the vortex onset-likely reflecting enhanced intermittency as flux elements are swept into the vortex-but later drops sharply between 1.5–3 Mm once the coherent structure is established. This entropy reduction suggests that the vortex acts to stabilize and organize , possibly through magnetic tension and vertical realignment. Prior to the onset of the vortex, both the magnitude and entropy of horizontal vorticity remain low in the upper atmosphere. This indicates a quiescent, nonrotational flow regime with minimal spatial variation, consistent with an irrotational, magnetically dominated background. The absence of pre-existing shear or turbulence suggests that the vortex did not arise from the gradual accumulation of small-scale swirling motions or steady twisting of field lines. Instead, it likely formed abruptly in response to a localized, possibly impulsive, event, potentially triggered by plasma or magnetic instabilities. The entropy of the kinetic energy () increases with vortex onset, decreasing immediately after, reflecting the emergence of coherent flow patterns. However, this trend reverses near the vortex’s half-life, where entropy increases significantly. This delayed rise is likely driven by the emergence of shear flows that act to destabilise the coherent vortex and inject kinetic complexity into the system. Notably, this entropy growth in is followed by a corresponding increase in the entropy of horizontal vorticity, which had been suppressed during the organised vortex phase. The subsequent increase in vorticity entropy marks the transition to a fragmented regime dominated by small-scale, disordered vortical structures, consistent with the formation of secondary swirls.
@@
Temperature and pressure both exhibit an initially high entropy in the lower atmosphere, reflecting significant spatial variability prior to vortex formation. As the vortex emerges, entropy in both quantities decreases. Notably, pressure entropy drops more rapidly than temperature entropy, indicating that the pressure field becomes ordered earlier, possibly due to the propagation of magnetoacoustic waves that redistribute pressure more efficiently than thermal energy. Following this initial organisation phase, the two variables diverge in behaviour. The pressure entropy remains low throughout the middle and upper atmosphere during the lifetime of the vortex, with only a modest increase near the footpoints and below 1 Mm. In contrast, temperature entropy begins to rise in the upper atmosphere as the vortex develops, pointing to the formation of fine-scale thermal structure likely driven by localised heating and dynamic mixing processes. This growing thermal complexity aligns with the evolution of plasma temperature shown in Figure 3, where the vortex core heats significantly during its mature phase, while pressure remains relatively stratified and stable in the same region.
@@
Figure 5 displays the evolution of the entropy in a control region without coherent vortex activity. A striking feature is the presence of vertically coherent, quasiperiodic fluctuations across several plasma variables, particularly in density, pressure, and temperature. These patterns likely reflect global p-mode oscillations driven by the lower boundary of the Bifrost simulation, which imposes perturbations to emulate solar acoustic modes. Interestingly, such periodic signatures are suppressed in the vortex region (Figure 4), suggesting that the presence of the vortex disrupts or decouples the propagation of these global oscillations. Notably, previous studies using the same simulation have identified magnetoacoustic waves propagating within the vortex core, particularly in the lower atmosphere. Thus, these waves, are likely generated or modulated by the vortex itself, rather than inherited from the bottom boundary.
2.4.1 Linear and nonlinear coupling
@@
To explore the degree of statistical dependence between rotational motion and key plasma parameters during the vortex lifecycle, we computed the NMI between the vorticity magnitude of horizontal plasma flow () and each relevant physical quantity as displayed in Figure 6 Here, MI values are normalized to fall between 0 (no statistical dependence) and 1 (perfect dependence). NMI in Figure 6 ranges from 0 to approximately 0.4, providing a dimensionless measure of statistical dependence between vertical vorticity and each plasma variable. Values close to 0 indicate little to no shared information, reflecting largely uncorrelated or independent behavior. Moderate NMI values ( 0.1–0.25) suggest partial coupling, where the variable exhibits some degree of organization influenced by the vorticity field, possibly through indirect or non-linear interactions. Higher NMI values (approaching 0.3–0.4) correspond to stronger statistical dependence, reflecting coordinated evolution. The panels reveal that, during the vortex’s lifetime, the coupling between vorticity and several thermodynamic and magnetic variables is locally enhanced, particularly in the upper atmosphere. This behaviour contrasts with the more uniform and subdued NMI patterns observed in the no-vortex control region (Figure 7), confirming that the enhanced coupling in the vortex core is dynamically driven rather than arising from static background structure. In other words, the elevated NMI between rotational flows and thermodynamic variables reflects genuine physical interaction associated with vortex dynamics, rather than a spurious correlation due to their co-location within the same magnetic flux tube.
@@
In the non–large-scale vortex region, where only small-scale swirling flows are present, NMI values remain generally low and spatially diffuse (Figure 7) in the upper atmosphere. This suggests a weak statistical coupling between vorticity and other plasma variables when coherent, vertically extended rotational structures are absent. The elevated NMI between temperature, , and vorticity in the photospheric control region likely arises from granular-scale convection, intergranular downflows, and p-mode-induced oscillations, all of which structure the plasma in ways that naturally correlate temperature, vertical motion, and rotational shear—even in the absence of a coherent vortex. Importantly, a low vorticity amplitude does not necessarily imply a low NMI, nor does a high vorticity guarantee strong coupling. What NMI captures is not the intensity of a given quantity, but the degree to which two variables co-vary in a structured and predictable manner. The persistently lower NMI in this control region reinforces the idea that strong mutual dependence, as seen in the vortex core, requires not just vorticity, but organized dynamics capable of modulating the surrounding plasma fields.

### v2

@@ 2.3.1 Quantifying statistical dependence with mutual information
In other words, knowing the state of one variable reduces the unpredictability of the other by that fraction. In practical terms, a normalized MI above 0.1 indicates a nontrivial statistical dependency. For MHD, this can reveal meaningful coupling between physical processes, such as energy transfer across fields, cross-scale interactions in turbulence, or correlations between magnetic and velocity structures that would not necessarily be visible in linear measures. Thus, even a seemingly “small” MI value may reflect deep physical connections within the plasma system, encompassing both dynamical and thermodynamic relationships. For two random variables and , their MI, , is defined as
@@
Before computing the NMI, each variable was rescaled to a fixed range, either or , depending on its physical bounds. Joint histograms were then constructed using 64 uniformly spaced bins across the normalized range. This fixed binning approach provides enough resolution to capture nonlinear dependencies while maintaining numerical stability in the mutual information estimates. Normalizing and binning in this way ensures consistency across variables and time steps, allowing for reliable comparisons and minimizing sensitivity to outliers or local fluctuations in the data. To assess the statistical reliability of the observed information-theoretic coupling, quantified by the NMI, we performed pixel-shuffle surrogate tests within each time-height window. For each driver-target variable pair (e.g., temperature as the driver and velocity or magnetic field components as the target), we generated spatial surrogates by randomly permuting pixel positions within the vortex mask. This procedure preserves the marginal amplitude distributions of each variable while destroying their spatial correspondence, thereby eliminating genuine statistical dependencies between them. Using the surrogate ensemble, we computed the MI values to construct a null distribution, from which the significance level, , was defined as the fraction of surrogate samples with MI values greater than or equal to the observed one. Across the entire analysed region and time interval, in nearly all pixels, with only isolated exceptions.
@@ Results
To characterize the temporal and spatial (in vertical direction) evolution of physical conditions within the vortex, we computed time-height maps of spatially averaged quantities inside the tracked magnetic flux tube core. By analyzing these averaged fields across the different vortex phases, before onset, development, and decay, we identify the signatures of dynamic coupling and energy concentration associated with vortex-driven activity. Figure 4 shows the temporal and height variation of the spatially averaged vorticity of horizontal velocity (), vertical velocity () and the base-10 logarithm of each of the other variables within the magnetic flux tube. The spatial averages were computed for each height plane. The stars of different colours indicate key stages in the vortex evolution as determined by the -method (Graftieaux2001; Giagkiozis_2018; Yuan2023), using the condition in the magnetic flux tube mask to define coherent rotation. We choose a lower value as initially the velocity field is not very twisted, leading to a lower value. Blue stars mark the onset of the main vortex at each height, while orange stars indicate the end of it. As the main vortex decays, it fragments into smaller secondary vortices that persist for a time before fully dissipating; their presence is marked by green stars. The vortex appears to form between the high photosphere (approximately 0.45 Mm) and the low chromosphere (around 1.0 Mm), after which it expands both upward into the corona and downward toward the solar surface. In the upper atmosphere, the vortex decays into multiple smaller structures, whereas in the lower layers it disappears more abruptly. In particular, smaller vortices are more persistent around 2 Mm, indicating that vortex evolution may be stratified with height, with different lifetimes and dynamics depending on atmospheric depth.
@@
The formation of the strong, coherent vortex leaves clear signatures in all plasma variables, especially in the chromosphere and transition region (above roughly 1 Mm). In the upper atmosphere, the vorticity changes sign around the time the vortex forms. This marks a shift from small, network-like flows that rotate clockwise to a larger, dominant vortex rotating counterclockwise. The positive vorticity linked to this large-scale vortex lasts through most of its lifetime but slowly weakens as the coherent rotation breaks apart. During the decay phase, the vortex transitions into smaller, fragmented swirling motions. This pattern shows that as the vortex develops, angular momentum spreads outward and energy shifts from the organized, large-scale rotation into smaller, irregular motions that can dissipate energy through heating.
@@
One of the most prominent features is a sharp and sustained drop in plasma-, which falls below unity shortly after vortex formation and remains low (log) as long as coherent vortical structures persist. This behaviour marks a transition from a gas-pressure-dominated to a magnetically dominated regime. Simultaneously, both the kinetic and magnetic energy increase by over an order of magnitude, peaking near vortex’s half-life. A local minimum in the vertical magnetic field coincides with the onset of the main vortex (blue stars). This may reflect either a redistribution of magnetic flux by the swirling plasma motions or the conversion of vertical field into horizontal components as the vortex twists and shears the local field. The subsequent increase in indicates magnetic flux accumulation within the vortex core as it matures.
@@
The formation of the vortex is accompanied by a coherent upflow of plasma, which transports denser fluid from lower atmospheric layers into the vortex core. This upflow results in a localized enhancement of mass density and a simultaneous drop in temperature, consistent with adiabatic expansion and the lifting of a cooler, denser plasma. Throughout the lifetime of the vortex, upward velocities dominate within the core region, continuously transporting plasma from lower layers and sustaining the elevated density. This persistent upflow plays a crucial role in maintaining the mass loading of the vortex structure. Only toward the end of the vortex’s life does this trend reverse, with downflows gradually becoming more prominent. This culminates in a strong downward jet along the vortex axis during its final decay phase, particularly in the upper atmospheric layers. The gas pressure initially increases at heights above 1.5 Mm as the upflow develops, but subsequently drops as the vortex matures. Following the initial cooling phase, the plasma within the vortex core is gradually heated, reaching temperatures up to an order of magnitude higher than the onset vortex state.
Figure 5 displays the evolution of the entropy in a control region without coherent vortex activity. All variables exhibit nearly periodic fluctuations resulting from global p-mode oscillations driven by the lower boundary of the Bifrost simulation, which imposes perturbations to emulate solar acoustic modes. The non-vortex region exhibits a more periodic and stratified behaviour, dominated by pressure-driven dynamics rather than magnetic stresses, indicating a relatively quiescent and stable plasma environment. It presents weaker magnetic fields and higher plasma values reflecting limited magnetic coupling and reduced energetic activity. A comparison with Figure 4 highlights fundamental differences in physical and dynamical properties. The vortex region lacks the clear periodic patterns seen in the non-vortex case, instead showing more irregular and complex temporal behaviour. It is characterized by stronger magnetic fields and enhanced magnetic twisting, particularly in the upper atmospheric layers, indicating that its dynamics are magnetically dominated. Correspondingly, the plasma is significantly lower in the vortex region, confirming that magnetic forces play a leading role in its evolution. The vortex region also exhibits a denser and progressively hotter atmosphere toward later times, consistent with magnetic energy concentration and conversion into thermal energy. In contrast, the non-vortex region maintains a more periodic, stratified, and plasma-pressure-dominated behaviour, typical of a less magnetically active environment.
@@ 2.4 Tracking Plasma Coupling and Coherence via Information Theory
The profiles presented in Figure 4 characterize the background evolution of the vortex core and provide a reference context to interpret subsequent information-theory and structural analyzes. The computed SE quantifies the local spatial variability of key physical quantities within the tracked magnetic flux tube where the solar tornado is developed. Figure 6 shows the temporal and vertical evolution of the SE for the same variables as in Figure 4, except for plasma-. The SE was computed within the dynamically tracked magnetic flux tube, and the entropy was estimated from the empirical probability distributions at each height and time step using a fixed binning scheme of 64 bins and logarithms in base 2, corresponding to a maximum possible entropy of bits. Each panel in Figure 6 presents the entropy values, revealing how the spatial organization and complexity of the plasma change over time and height. In practice, entropy values near – correspond to relatively ordered fields with low spatial variability, while intermediate values (–) indicate partially structured states where coherent features coexist with moderate fluctuations. Higher values approaching reflect strongly intermittent fields, likely associated with turbulence, sharp spatial gradients, or enhanced small-scale variability.
@@
The onset and development of the main vortex are marked by significant changes in the spatial organization of magnetic fields, as captured by entropy diagnostics. During the formation phase, the entropy of magnetic energy () rises in the lower atmosphere, indicating increasing spatial complexity and localized structuring of the field. This is accompanied by a divergence between the entropy of and that of magnetic shear (), consistent with the formation of current sheets rather than uniform shear enhancement. While becomes more intermittent due to twisting and compression by the vortex, magnetic shear entropy remains vertically patchy and fragmented, suggesting that complexity arises from thin, spatially confined features rather than widespread braiding. Similarly, the entropy of the vertical field component () increases during the vortex onset-likely reflecting enhanced intermittency as flux elements are swept into the vortex-but later drops sharply between 1.5-3 Mm once the coherent structure is established. This entropy reduction suggests that the vortex acts to stabilize and organize , possibly through magnetic tension and vertical realignment. Before the onset of the vortex, both the magnitude and entropy of vorticity of horizontal plasma flows remain low in the upper atmosphere. This indicates a quiescent, non-rotational flow regime with minimal spatial variation, consistent with an irrotational, magnetically dominated background. The absence of pre-existing shear or turbulence suggests that the vortex did not arise from the gradual accumulation of small-scale swirling motions or steady twisting of field lines. Instead, it likely formed abruptly in response to a localized, possibly impulsive, event, potentially triggered by plasma or magnetic instabilities. The entropy of the kinetic energy () increases with vortex onset, decreasing immediately after, reflecting the emergence of coherent flow patterns. However, this trend reverses near the vortex’s half-life, where entropy increases significantly. This delayed rise is likely driven by the emergence of shear flows that act to destabilise the coherent vortex and inject kinetic complexity into the system. Notably, this entropy growth in is followed by a corresponding increase in the entropy of vorticity, which had been suppressed during the organized vortex phase. The subsequent increase in vorticity entropy marks the transition to a fragmented regime dominated by small-scale, disordered vortical structures, consistent with the formation of secondary swirls.
@@
Temperature and pressure both exhibit an initially high entropy in the lower atmosphere, reflecting significant spatial variability before the vortex formation. As the vortex emerges, entropy in both quantities decreases. Notably, pressure entropy drops more rapidly than temperature entropy, indicating that the pressure field becomes ordered earlier, possibly due to the propagation of magnetoacoustic waves that redistribute pressure more efficiently than thermal energy. Following this initial organisation phase, the two variables diverge in behaviour. The pressure entropy remains low throughout the middle and upper atmosphere during the lifetime of the vortex, with only a modest increase near the footpoints and below 1 Mm. In contrast, temperature entropy begins to rise in the upper atmosphere as the vortex develops, pointing to the formation of fine-scale thermal structure likely driven by localised heating and dynamic mixing processes. This growing thermal complexity aligns with the evolution of plasma temperature shown in Figure 4, where the vortex core heats significantly during its mature phase, while pressure remains relatively stratified and stable in the same region.
@@
Figure 7 displays the evolution of the entropy in a control region without coherent vortex activity. A striking feature is the presence of vertically coherent, nearly periodic fluctuations across several plasma variables, particularly in density, pressure, and temperature. These patterns likely reflect global p-mode oscillations driven by the lower boundary of the Bifrost simulation, which imposes perturbations to emulate solar acoustic modes. These acoustic modes periodically compress and expand the plasma, modulating temperature, density, and pressure gradients. As a result, the spatial complexity of these quantities, measured through SE, oscillates in phase with the p-mode cycles. Thus, higher SE values correspond to the compressive phases with steeper gradients, and lower values to the rarefactive, smoother phases. Interestingly, such periodic signatures are suppressed in the vortex region (Figure 6), suggesting that the presence of the vortex alters or even decouples the propagation of global p-mode oscillations. The strong and twisted magnetic fields within the vortex modify the local wave environment, favouring the generation and propagation of magnetoacoustic modes confined within the magnetic structure. Previous analyses of the same simulation have identified such waves in the lower atmosphere, supporting the idea that the oscillatory behaviour within the vortex is primarily driven or modulated internally, by the vortex dynamics themselves, rather than directly inherited from the p-mode forcing at the lower boundary.
2.4.1 Flow-Rotation Coupling with Plasma Parameters
@@
To explore the degree of statistical dependence between rotational motion and key plasma parameters during the vortex lifecycle, we computed the NMI between the vorticity magnitude of horizontal plasma flow () and each relevant physical quantity as displayed in Figure 8. Here, MI values are normalized, so that MI reflects no statistical dependence, while MI denotes perfect dependence. NMI in Figure 8 ranges from 0 to approximately 0.4, providing a dimensionless measure of statistical dependence between vertical vorticity and each plasma variable. Values close to 0 indicate little to no shared information, reflecting largely uncorrelated or independent behavior. Moderate NMI values ( 0.1-0.3) suggest partial coupling, where the variable exhibits some degree of organization influenced by the vorticity field, possibly through indirect or non-linear interactions. Higher NMI values (0.3-0.4) indicate a stronger statistical dependence between variables, consistent with coordinated evolution. During the vortex phase, the coupling between vorticity and several thermodynamic and magnetic quantities increases markedly, particularly in the upper atmosphere. The control region without a vortex exhibits only weak and spatially uniform NMI patterns (Figure 9). This contrast shows that the enhanced coupling within the vortex core is a dynamic signature of vortex activity, not a static property of the background magnetic field.
@@
In the non-large-scale vortex region, where only small-scale swirling flows are present, NMI values remain generally low and spatially diffuse (Figure 9) in the upper atmosphere. This suggests a weak statistical coupling between vorticity and other plasma variables when coherent, vertically extended rotational structures are absent. The elevated NMI between temperature, , and vorticity in the photospheric control region likely arises from granular-scale convection, intergranular downflows, and p-mode-induced oscillations, all of which structure the plasma in ways that naturally correlate temperature, vertical motion, and rotational shear—even in the absence of a coherent vortex. Importantly, a low vorticity amplitude does not necessarily imply a low NMI, nor does a high vorticity guarantee strong coupling. What NMI captures is not the intensity of a given quantity, but the degree to which two variables co-vary in a structured and predictable manner. The persistently lower NMI in this control region reinforces the idea that strong mutual dependence, as seen in the vortex core, requires not just vorticity, but organized dynamics capable of modulating the surrounding plasma fields.